If a, b, c are real numbers and satisfy 1/a + 1/b + 1/c = 1/(a+b+c), then the value of ((1/a + 1/b + 1/c)^2007) - (2007 / ( a^2007 + b^2007 + c^2007 )) + ( 2006 / ((a + b + c)^2007 )) is :
(a)2
(b)0
(c)3
(d)1

Grade:12

3 Answers

Anoopam Mishra
9 years ago
I don’t the proper method to solve this but you can solve it by this method.
Take a=1, b=-1, c=1, then above equation is also satisfied.
Value of expression = 1-2007+2006 = 0
So the answer is (b).
Anoopam Mishra
9 years ago
Please approve.
manmath
8 years ago
answer is b
 
I don’t the proper method to solve this but you can solve it by this method.
Take a=1, b=-1, c=1, then above equation is also satisfied.
Value of expression = 1-2007+2006 = 0
So the answer is (b).